OpenStudy (anonymous):

how do i solve x and y? -2x+7y=22 x=7y+3 please help me :(

OpenStudy (anonymous):

want me to solve it for you

OpenStudy (anonymous):

or steps

OpenStudy (anonymous):

please????

OpenStudy (anonymous):

solve???

OpenStudy (raden):

see the 2nd equation : x=7y+3 replace the 7y to left side, it becomes x - 7y = 3 now add both equations above : -2x + 7y = 22 x - 7y = 3 ------------- + -x = 25 x = -25 to get y, subtitute x = -25 into the 1st equation. -2x + 7y = 22 -2(-25) + 7y = 22 50 + 7y = 22 7y = 22 - 50 7y = -28 y = -28/7 y = -4

OpenStudy (raden):

hoped that make sense for you, @LaBlueEyes2013